# Predictions of the Nancy Grace Roman Space Telescope Galactic Exoplanet Survey. IV. Lens Mass and Distance Measurements

### v1

@@ 3.2 White Dwarf & Very Low Mass Stellar Hosts
We expect ${\sim}20$% of all lens hosts to be comprised of white dwarfs (WDs) or brown dwarfs (BDs) (Gould, 2000; Bennett et al., 2007; Gaudi, 2012), with a $\lesssim 1\%$ contribution from neutron stars and black holes, which we ignore here. Along with WD and BDs, very low mass (VLM) at the bottom of the main sequence are also expected to make up a small fraction of the lens hosts, these objects are also very dim. We note the isochrones used to generate the stellar population in GULLS have a lower-limit of $0.1M_{\odot}$, which is $0.02M_{\odot}$ above the lower mass limit of the stellar population that has been generated via SynthPop. Therefore we denote the range of host masses between $0.08-0.10M_{\odot}$ as VLMs. These hosts do not have a measurable brightness in any Roman passband, and we group them into the dark category along with WD hosts. Currently GULLS does not support BDs or BD isochrones, so our subsequent analysis does not include planet-hosting BDs.
As a check on our sampled population, we confirm that WDs and VLMs make up an appropriate proportion (in our case $\sim$}19%) of the 3,000 events we’ve selected for processing with pyLIMASS. Figures 3, 4, and 5 distinguish the sample of stellar hosts (e.g. luminous lenses) from the sample of dark hosts (non-luminous lenses). Since all of these dark objects do not have appreciable flux in Roman, a precise determination of the lens mass and distance must rely on both higher-order effects to be well-measured from the simulated light curve photometry, as previously explained in Section 2.1. Lastly, it may be possible to estimate $thete_{\rm E}$ via precise astrometry of the source, whose centroid shifts during the microlensing event due to the time-varying fluxes and positions of the unresolved micro-images created during the event. This signature would be largest for the most massive and closest host lenses, and may offer an avenue to improve the fraction of events with constraints on the lens masses and distances (Dominik & Sahu, 2000; Lam et al., 2023; Sajadian & Sahu, 2023). This is a topic of active investigation within the RGES PIT, however it is beyond the scope of work in our current study.
@@
Bachelet et al. (2024) simulated 1000 PSPL microlensing events observed in Roman passbands F087 and F146, including details of the observing strategy given in Penny et al. (2019) (50s exposure time, zero-point magnitude of $Z_{P}=27.4$ mag). The observables that were used for these simulated events are: $t_{E},F087_{S},F146_{S},A_{V_{S}},F087_{L},F146_{L},\theta_{*},\pi_{\textrm{EN}},\pi_{\textrm{EE}},\mu_{\textrm{rel}}$. Figure 3 of Bachelet et al. (2024) shows the distributions of the accuracies with which the lens masses and distances of these simulated events are reconstructed, specifically the distributions of $\delta M_{L}$ and $\Delta D_{L}$, the difference between the values of $M_{L}$ and $D_{L}$ inferred from pyLIMASS and their true values. A correlation can also be seen in their distribution of source-lens brightness difference ($\Delta F146$) as a function of lens mass. This is because $\theta_{E}$ can be measured from the ingested observables ($t_{E}$ and $\mu_{\textrm{rel}})$, and the lens mass can be approximated by $M_{L}\,{\approx}\,\theta_{E}^{2}D_{L}/\kappa$, where $\kappa=4G/c^{2}$. They demonstrated that the medians of the $\delta M_{L}$ and $\Delta D_{L}$ values are close to zero, although they do find a bias in $M_{L}$ for events where the source and/or lens brightness are poorly measured.
@@ 4.1.1 Source Distance Prior via Galactic Models
When attempting to determine the distance to the lens system, one must take into account the distance to the source via:
$D_{L}=\frac{\textrm{AU}}{\pi_{E}\theta_{E}+\textrm{AU}/D_{S}}$ (11)
@@
It is worth noting that there is a difference in the isochrones used to generate the stellar population in SynthPop (MIST), and the isochrones that pyLIMASS uses for its mass-luminosity relations (PARSEC). This means that our analysis is effectively probing the systematic uncertainty due to the different choices of isochrones. Therefore, the magnitude of the differences between the isochrones represents the systematic uncertainties in our ability to model the true mass-luminosity relationship of the stars of interest. It would appear that this systematic uncertainty in our inferred parameters due to this source of astrophysical uncertainty is relatively small given our results presented in Section 4.2.
@@
Finally, we use pyLIMASS to determine lens masses and distances and associated uncertainties for all 3000 events in our final sample. For the input observables, we have included light curve fitting parameters with associated fisher uncertainties from Zohrabi et al. (in prep) (Figure 2), as well as estimates of the lens flux (and errors) and estimates of lens-source relative proper motion (and errors) via the image elongation and color-dependent centroid shift methods (Verma et al., in prep) (Section 2.3). For the ${\sim}20\%$ of simulated events with dark lens hosts, there is no measurement of lens flux or lens-source relative proper motion, so those observables have been omitted.
@@
Figure 5 shows the comparisons of $M_{L}$ (left panel) and $D_{L}$ (right panel) for the true values relative to the pyLIMASS-estimated values (and uncertainties). The inferred values of $M_{L}$ and $D_{L}$ for the stellar (luminous) population agrees well with the true values. The dark population of lenses are in poorer agreement, for the reasons discussed earlier. We note that there is a peak in the distance distribution of lenses near $D_{L}\sim 7.5~{\rm kpc}$, implying that the majority of exoplanet systems that Roman will detect may be in the bulge, although of course this depends on how the frequency of exoplanets depends on host star properties and environment.
@@ 6 Discussion and Conclusion
Figure 7: A representative sample of nine historical events from Table 3. The top row shows nearby (<4<4 kpc) events for which a combination of all three mass-distance relations ($\theta_{E},\pi_{E}$, lens flux) give an unambiguously measured lens mass and distance. The yellow data with error bars shows the best fit result from image-constrained modeling, the blue contours give the posterior distribution (central 39%, 86%, 99%) from pyLIMASS. The middle row shows more distant events (>4>4 kpc) for which the estimate of lens mass and distance is still relatively accurate but less precise. The bottom row shows problematic events for which there are degeneracies that remain (MB10328; Vandorou et al. 2025b), or the identification of the lens is dubious (MB11262; Terry et al. 2025b), or the significantly overlapping $\theta_{E}$ and $\pi_{E}$ relations result in large errors (OB141760; Rektsini et al. 2025).
@@ Acknowledgments
Funding for the Roman Galactic Exoplanet Survey Project Infrastructure Team is provided by the Nancy Grace Roman Space Telescope Project through the National Aeronautics and Space Administration grant 80NSSC24M0022, by The Ohio State University through the Thomas Jefferson Chair for Space Exploration endowment, and by the Vanderbilt Initiative in Data-intensive Astrophysics (VIDA). JPB has been supported by the Australian Government through the Australian Research Council Discovery Project Grants 240101842, JPB and AC are supported by the SPACE-MLENS ANR grant ANR-24-CE31-3263.

### v2

@@ 3.2 White Dwarf & Very Low Mass Stellar Hosts
We expect ${\sim}20$% of all lens hosts to be comprised of white dwarfs (WDs) or brown dwarfs (BDs) (Gould, 2000b; Bennett et al., 2007; Gaudi, 2012), with a $\lesssim 1\%$ contribution from neutron stars and black holes, which we ignore here. Along with WD and BDs, very low mass (VLM) at the bottom of the main sequence are also expected to make up a small fraction of the lens hosts, these objects are also very dim. We note the isochrones used to generate the stellar population in GULLS have a lower-limit of $0.1M_{\odot}$, which is $0.02M_{\odot}$ above the lower mass limit of the stellar population that has been generated via SynthPop. Therefore we denote the range of host masses between $0.08-0.10M_{\odot}$ as VLMs. These hosts do not have a measurable brightness in any Roman passband, and we group them into the dark category along with WD hosts. Currently GULLS does not support BDs or BD isochrones, so our subsequent analysis does not include planet-hosting BDs.
As a check on our sampled population, we confirm that WDs and VLMs make up an appropriate proportion (in our case $\sim$}19%) of the 3,000 events we’ve selected for processing with pyLIMASS. Figures 3, 4, and 6 distinguish the sample of stellar hosts (e.g. luminous lenses) from the sample of dark hosts (non-luminous lenses). Since all of these dark objects do not have appreciable flux in Roman, a precise determination of the lens mass and distance must rely on both higher-order effects to be well-measured from the simulated light curve photometry, as previously explained in Section 2.1. Lastly, it may be possible to estimate $thete_{\rm E}$ via precise astrometry of the source, whose centroid shifts during the microlensing event due to the time-varying fluxes and positions of the unresolved micro-images created during the event. This signature would be largest for the most massive and closest host lenses, and may offer an avenue to improve the fraction of events with constraints on the lens masses and distances (Dominik & Sahu, 2000; Lam et al., 2023; Sajadian & Sahu, 2023). This is a topic of active investigation within the RGES PIT, however it is beyond the scope of work in our current study.
@@
Bachelet et al. (2024) simulated 1000 PSPL microlensing events observed in Roman passbands F087 and F146, including details of the observing strategy given in Penny et al. (2019) (50 sec exposure time, zero-point magnitude of $Z_{P}=27.4$ mag). The observables that were used for these simulated events are: $t_{E},F087_{S},F146_{S},A_{V_{S}},F087_{L},F146_{L},\theta_{*},\pi_{\textrm{EN}},\pi_{\textrm{EE}},\mu_{\textrm{rel}}$. Figure 3 of Bachelet et al. (2024) shows the distributions of the accuracies with which the lens masses and distances of these simulated events are reconstructed, specifically the distributions of $\delta M_{L}$ and $\Delta D_{L}$, the difference between the values of $M_{L}$ and $D_{L}$ inferred from pyLIMASS and their true values. A correlation can also be seen in their distribution of source-lens brightness difference ($\Delta F146$) as a function of lens mass. This is because $\theta_{E}$ can be measured from the ingested observables ($t_{E}$ and $\mu_{\textrm{rel}})$, and the lens mass can be approximated by $M_{L}\,{\approx}\,\theta_{E}^{2}D_{L}/\kappa$, where $\kappa=4G/c^{2}$. They demonstrated that the medians of the $\delta M_{L}$ and $\Delta D_{L}$ values are close to zero, although they do find a bias in $M_{L}$ for events where the source and/or lens brightness are poorly measured.
@@ 4.1.1 Source Distance Prior via Galactic Models
Recall that a measurement of the distance to the lens through measurements of $\theta_{E}$ and $\pi_{E}$ depends on the distance to the source via Equation LABEL:eqn:dl-sl.
@@
A subtle but important caveat worth mentioning is that there is a difference in the isochrones used to generate the stellar population in SynthPop (MIST), and the isochrones that pyLIMASS uses for its mass-luminosity relations (PARSEC). This means that our analysis is effectively probing the systematic uncertainty due to the different choices of isochrones. Therefore, the magnitude of the differences between the isochrones represents the systematic uncertainties in our ability to model the true mass-luminosity relationship of the stars of interest. It would appear that this systematic uncertainty in our inferred parameters due to this source of astrophysical uncertainty is relatively small given our results presented in Section 4.2.
@@
Finally, we use pyLIMASS to determine lens masses, distances, and associated uncertainties for all 3000 simulated events in our final sample. For the input observables, we have included light curve fitting parameters with associated fisher uncertainties from Zohrabi et al. (in prep) (Figure 2), as well as estimates of the lens flux (and errors) and estimates of lens-source relative proper motion (and errors) via the image elongation and color-dependent centroid shift methods (Verma et al., in prep) (Section 2.3). For the ${\sim}20\%$ of simulated events with dark lens hosts, there is no measurement of lens flux or lens-source relative proper motion, so those observables have been omitted.
Figure 5: Estimated masses and distances for a representative sample of simulated events. The constraints from an estimate of $\theta_{E}$ are shown as brown shaded curves, and the $\pi_{E}$ constraints are shown as teal shaded curves. The blue, green, and red lines show the constraint from the measured flux in the Roman $F087$, $F146$, and $F213$ passbands, combined with mass-luminosity relations from the pyLIMASS-implemented PARSEC isochrones (Bressan et al., 2012). The dashed lines show the $1\sigma$ errors that come from the quadrature sum of the lens flux uncertainties and an assumed 5% intrinsic uncertainty on the mass-luminosity relations. The yellow data points show the true value of the lens mass and distance. The blue contours give the pyLIMASS posterior distributions (central 39%, 86%, 99%) for stellar (luminous) lenses, red contours (bottom row) give the posteriors for three dark lenses (two WDs and one VLM star).
@@
We have selected a sub-sample of nine simulated events to display in Figure 5. The events are evenly distributed across the GBTDS footprint and we have sampled at least one event from each planet mass bin. Each panel shows a wide lens mass and distance parameter space ($0.01M_{\odot}<M_{L}<1.0M_{\odot}$ and $0.5$ kpc $D_{L}<9$ kpc). The top two rows of Figure 5 show stellar (luminous) lenses, and the bottom row shows dark lenses. The combination of lens flux detections and at least one of two light curve higher-order effects will enable robust lens mass and distance estimates during the GBTDS. In particular, this figure demonstrates the importance of lens flux measurements in (most) cases where the microlensing parallax is not well-measured from the light curve data.
Figure 6 shows the comparisons of $M_{L}$ (left panel) and $D_{L}$ (right panel) for the true values relative to the pyLIMASS-estimated values (and uncertainties) for the simulated events in the sample. The inferred values of $M_{L}$ and $D_{L}$ for the stellar (luminous) population agree well with the true values. The dark population of lenses are in poorer agreement, for the reasons discussed earlier. We note that there is a peak in the distance distribution of lenses near $D_{L}\sim 7.5~{\rm kpc}$, implying that the majority of exoplanet systems that Roman will detect may be in the bulge, although of course this depends on how the frequency of exoplanets depends on host star properties and environment. For clarity in Figure 6 we remove outliers with pyLIMASS-estimated $M_{L}>3\sigma$ from the true $M_{L}$ value, this accounts for ${\sim}6$% of simulated events. This population of outliers may be due to expected statistical fluctuations, subtle biases inherent in pyLIMASS, true degeneracies in some of the simulated events, or likely a combination of these factors. We note that less than $0.5\%$ of simulated events are $5\sigma$ outliers in $M_{L}$.
@@ 6 Discussion and Conclusion
Figure 8: Similar to Figure 5, but for a representative sample of nine historical events from Table 3. The top row shows nearby (<4<4 kpc) events for which a combination of all three mass-distance relations gives an unambiguously measured lens mass and distance. The constraints from $\theta_{E}$, $\pi_{E}$, and lens flux are shown as the brown, teal, and red/blue/green curves. The lens flux constraints in these cases come from empirical $V$, II, and $K$ band mass-luminosity relations of Delfosse et al. (2000). The yellow data with error bar shows the best fit result from image-constrained modeling, and the blue contours give the posterior distribution (central 39%, 86%, 99%) from pyLIMASS. The middle row shows more distant events (>4>4 kpc) for which the estimate of lens mass and distance is still relatively accurate but less precise. The bottom row shows problematic events for which there are degeneracies that remain (MB10328; Vandorou et al. 2025b), the identification of the lens is dubious (MB11262; Terry et al. 2025b), or the significantly overlapping $\theta_{E}$ and $\pi_{E}$ relations result in large errors (OB141760; Rektsini et al. 2025).
@@ Acknowledgments
Funding for the Roman Galactic Exoplanet Survey Project Infrastructure Team is provided by the Nancy Grace Roman Space Telescope Project through the National Aeronautics and Space Administration grant 80NSSC24M0022, by The Ohio State University through the Thomas Jefferson Chair for Space Exploration endowment, and by the Vanderbilt Initiative in Data-intensive Astrophysics (VIDA). JPB has been supported by the Australian Government through the Australian Research Council Discovery Project Grants 240101842, JPB and AC are supported by the SPACE-MLENS ANR grant ANR-24-CE31-3263. MJH Acknowledges support from the Heising-Simons Foundation under grant No. 2022-3542.